OpenStudy (anonymous):

Hillary changes the oil in her car after every 6,000 kilometers. If Hillary drives 970 kilometers each month, about how many months will pass between each oil change?

OpenStudy (anonymous):

plz. help

jimthompson5910 (jim_thompson5910):

Let x = number of months

jimthompson5910 (jim_thompson5910):

In one month, she drives 970*1 = 970 km In 2 months, she drives 970*2 = 1940 km In 3 months, she drives 970*3 = 2910 km ... ... So in x months, she drives 970x kilometers

jimthompson5910 (jim_thompson5910):

Simply set this equal to 6000 and solve for x

jimthompson5910 (jim_thompson5910):

So you need to solve 970x = 6000

OpenStudy (anonymous):

it is 6 months

jimthompson5910 (jim_thompson5910):

Roughly 6.1856 months, so if you round down, then yes...it's roughly 6 months

OpenStudy (anonymous):

thanks so much

jimthompson5910 (jim_thompson5910):

yw